OpenStudy (kj4uts):

Given f(x) = x / (x^2 - 4), find and simplify f (3x).

OpenStudy (kj4uts):

Are they asking me to plug 3(x) into the equation like this 3x / (3x^2 - 4)? If that is the case I do not think it can be simplified anymore.

OpenStudy (xmissalycatx):

Let's see..

OpenStudy (kittiwitti1):

\[f(x)=\frac{x}{x^{2}-4}\rightarrow f(3x)=\frac{3x}{(3x)^{2}-4}=\frac{3x}{9x^{2}-4}\]

OpenStudy (kj4uts):

Oh I see so you square the 3^2 to get the 9 on the bottom?

OpenStudy (kittiwitti1):

\[\rightarrow\frac{3x}{9x^{2}-4}\rightarrow\frac{3x}{(3x-2)(3x+2)}\]

OpenStudy (xmissalycatx):

I was drawing it out, but every time I'm about to send what I'm drawing @kittiwitti1 sends those equations XD poop

OpenStudy (kittiwitti1):

Well... if you don't put it in the equation properly you end up getting\[3x^{2}\text{ which is basically like }3\times x^{2}\]But what we're looking for here is\[(3x)^{2}\text{ as you're putting } "3x"\text{ into the place of }x\]LOL sorry xD

OpenStudy (kittiwitti1):

It's asking you to simplify, you don't need to reverse the process and re-multiply... I just factored it out for you. Lol

OpenStudy (kj4uts):

Thank you for your time and help :)

OpenStudy (kittiwitti1):

No problem ^-^

OpenStudy (kittiwitti1):

By the way, if you get confused there is always WolframAlpha to check your problems with~ http://www.wolframalpha.com/input/?i=f(x)+%3D+x+%2F+(x%5E2+-+4)+when+x%3D3x